1. A 40-kg person is riding on a Ferris wheel of r
Answer #1

m=40kg

r= 20 m

v= 6m/s

1. centripetal acceleration = v2/r =36/20 = 1.8 m/s2

2. At bottom, net force acting = mg + mv2/r = 40 x 9.8 + 40 x 1.8 = 464 N

At top, net force acting = mg - mv2/r = 40 x 9.8 - 40 x 1.8 = 320 N

3. m1 = m2 =m = 2 x 1030 kg

r= 3 x 1012 m

F= Gmm/r2

= (6.67 x 10-11 ) ( 2 x 1030) ( 2 x 1030)/(3 x 1012)2

= 1.48 x 10-5 N